OpenStudy (aurorab):

2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 (3x+6 8 ) = 24 1 (7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

That equation is really hard to read, but to check the answer, all Max would have to do is plug what he got for X back into the original equation.

OpenStudy (aurorab):

3x+6 = 7x−1 8 6 24 (3x+6) = 24 (7x−1) 1 8 1 6 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (aurorab):

is it more understandable now?? \:

OpenStudy (anonymous):

What are the numbers under your equations?

OpenStudy (aurorab):

there supposed to be fractions... 24/1 and so forth

OpenStudy (anonymous):

Oh, yeah. I still can't understand that. Can you use the / or the equation button?

OpenStudy (aurorab):

yea..?

OpenStudy (anonymous):

The answer to this question is wrong because if you read the problem carefully, it said that Max wasn't sure if his answer is correct. He can check his answer by plugging in his x value, after he fixes his multiplication error in step 3. If 4 *7x is 28, then 4*-1 is not -1. The answer would be -4. Therefore step 3 would read as followed: 9x+18=28x-4 rather than: 9x+18-28x-1. The correct solution and how to check if its correct is as followed: 3(3x+6) = 4(7x-1) 9x+18 = 28x-4 18 = 19x-4 22 = 19x 22/19 = x How to check the answer: 3(22/19+6)/8 = 7(22/19-1)/6 45/38 = 45/38 ( Is true)